# Activation Addition: Steering Language Models Without Optimization

### v4

@@ 4 Results
A summary of all experiments can be found in Table 5.
Natural-language control of sentiment, topic, style. See Figure 2 and Tables 3 and 4 for benchmark results. We report some illustrative results in Table 1 and Appendix Table 6, including vectors for inducing anger (sentiment steering), weddings (topic steering), and conspiracy theories (topic/style steering). These completions were all obtained through one run of top-3 sampling. In particular, note the Eiffel vector (fact editing reminiscent of the ROME method (Meng et al. 2023), the range of sentiments and styles induced, and (to a weaker extent) the alignment (see the Hurt vector in Appendix Table 6). A runnable notebook with all examples is at tinyurl.com/actadd3.
@@
Intervention effectiveness. To evaluate how effectively ActAdd  steers output, we test on the OpenWebText corpus (Peterson et al. 2018). Our running example is the “wedding” topic vector produced by setting $p_{+}=$ weddings, $p_{-}=$    , $l=16$, $c=1$). See ‘Generality’ for broader tests.
@@
Experiment 1: perplexity ratio. First we check that ActAdd  increases the probability of the model outputting tokens related to the wedding vector. Appendix Figure 6 shows the resulting perplexity ratios under ActAdd, relative to the unmodified model. On documents where the injected topic is more relevant, ActAdd’s relative predictive performance increases.
@@
Experiment 2: logprob distribution shift. However, is the intervention affecting the tokens we expect it to, or is it reducing perplexity in some spurious way? Appendix Figure 9 shows the resulting mean log-probability difference distribution. We see that is approximately normal for the bulk of the tokens but with clearly heavy tails. The positive tail is significantly heavier than the negative tail, suggesting that one set of tokens are reliably increased in probability, with a smaller set of tokens reliably decreased to a lesser extent. Outlier tokens can be found in Appendix Table 11. From this we see clearly that the probabilities most increased on average are primarily wedding-related. The bottom tokens share no obvious theme and show a significantly lower absolute change in probability.
@@
Experiment 3: generation-scoring. Appendix Figures 7 and 8 show the result of sweeping the ActAdd intervention over layers and taking the count or fraction of related words. We see that the intervention (in this case) is already effective at the very first layer, rises in effectiveness until $l=6$, and then declines. Note also that for the optimal injection site we see $>$90% success in steering topic (compared to a $\sim$2baseline).
Generality. The wedding steering vector used above is just a running example to keep the impact concrete. (Averaging over a collection of steering vectors makes it difficult to depict the magnitude of the steering.) To show this topic is not a special case, we test a range of topics against a generic prompt, and use GPT-3.5 to score relative relevance (i.e. whether the steering worked). The results can be seen in Figure 2: we see a large effect (c. 5-20% on GPT’s gestalt measure) on all topics at strength $c=2$, with the exception of ‘art’.
@@
Toxicity reduction. To establish a common scale, we reused the baselines and PREADD results from Pei et al. 2023, adding results from Zhong et al. 2023. This yields 6 baselines to compare ActAdd against. We also considered Gu et al. 2022 (which reported 0.043 toxicity), but we could not reproduce the results; also, their disfluency (54.6) is too high for practical use. We compare to ActAdd using OPT (Zhang et al. 2022b) and LLaMA-3 (Meta 2024). Table 3 (matching methods on pretrained models) showed ActAdd-OPT has 8% lower toxicity than the second-best, PREADD-D-OPT, and ActAdd-LLaMA-3 gives a 5% drop over LLaMA-3 with a very small fluency penalty.
Sentiment control. Table 4 shows that our method is competitive on a conventional measure of sentiment control (Maas et al. 2011). We obtain state of the art success at steering from negative to positive sentiment; and the only method which outperforms ActAdd in the positive to negative direction incurs a large penalty to fluency (68.4 vs 24.2, when matching methods on the same pretrained model) and relevance. We speculate that the difficulty of steering LLaMa-3 towards negative sentiment is due to the output alignment methods used in its post-training. See Appendix E for details.
Table 3: Results on RealToxicityPrompts (random n=1000). The OPT used is 6.7B parameters, LLaMA-3-8B. Bold is $p<0.05$ against second-best. Gray text denotes numbers reported by Pei et al. 2023 (PREADD), Zhong et al. 2023 (Air-Decoding). More recent models are less toxic by default, but, even so, ActAdd steering an OPT model outperforms an unsteered LLaMA-3.
Control Type Method Toxicity $\downarrow$ (Dis)Fluency $\downarrow$ Relevance $\uparrow$
Steering vector ActAdd-OPT .112 13.8 .329
Steering vector ActAdd-LLaMA3 .108 6.7 .365
@@
Prompting OPT .200 54.3 .294
Controlled generation FUDGE (k = 100) .128 22.1 .329
Contrast decoding PREADD-S-OPT .134 51.7 .290
Contrast decoding PREADD-D-OPT .122 56.6 .326
Gradient-guided gen Air-Decoding .185 48.3 -
Table 4: Results on IMDb sentiment (random n=1000). Suc $=$ success, probability of changing sentiment classification. Flu $=$ (dis)fluency; perplexity. Rel $=$ relevance. We follow the convention of using ‘Fluency’ to mean disfluency (lower better). OPT is 6.7B, LLaMA-3 is 8B parameters. Bold results represent $p<0.05$ compared to the second-best. Gray text denotes numbers reported by Pei et al. 2023. Underline denotes best steered result. Fluency is worse under all steering methods; 1.5x to 3x worse for ActAdd, 7x worse for PREADD.
@@
Control Method Suc $\uparrow$ Flu $\downarrow$ Rel $\uparrow$ Suc $\uparrow$ Flu $\downarrow$ Rel $\uparrow$
Steering vector ActAdd-OPT 0.432 24.2 0.387 0.564 20.95 0.363
Steering vector ActAdd-LLaMA3 0.268 8.6 0.354 0.669 15.2 0.275
Unsteered OPT 0.175 8.95 0.430 0.445 9.38 0.423
Unsteered LLaMA3 0.138 5.8 0.437 0.417 6.09 0.426
Prompting OPT 0.307 53.5 0.298 0.365 50.9 0.287
Controlled gen FUDGE k$=$100 0.532 25.1 0.311 0.551 22.7 0.320
Contrast decoding PREADD-S-OPT 0.631 68.4 0.253 0.624 67.1 0.258
Preservation of off-target performance. Per Figure 3, on the ConceptNet benchmark of factual questions, our method has a negligible impact on off-target answer probabilities.
@@
[…]
Algebraic combination of forward passes. Every example of ActAdd  in this paper (e.g. those in Appendix Table 6) is an example of composing forward passes (e.g. we compose $\mathbf{h}_{+}$, $\mathbf{h}_{-}$ and $\mathbf{h}^{*}$ to produce steered output). This forms evidence for compositional representations (Olah 2023), independent of the evidence from task-composition arithmetic on weights (Ilharco et al. 2023). See Appendices D and E for further experiments into the mechanism of a steering vector.
Scales with model size. We estimate the relationship between ActAdd inference overhead and model size, to check that the method will remain relevant for massive frontier models and future models. Because ActAdd  involves only forward passes, it scales naturally with model size (Figure 4): the relationship between inference time premium and model size is decreasing.
@@ 5 Discussion
ActAdd  works well in some cases, but it still requires a search for its $p_{+},c$ and $l$ arguments. (So far we have had success with fixing $a=1$.) This makes it less user-friendly than simple prompt engineering. We include examples of failed steering vectors in Appendix Table 7; in particular, we lack understanding of when large injection coefficients damage capabilities. Further, even in its 1.5B form, GPT-2  is not sophisticated enough to support any demonstration of the effect of ActAdd  on reasoning tasks. ActAdd  still has two free parameters, the injection coefficient and the target layer, and if these values were very sensitive to the desired contrast pair then the method would suffer some computational overhead. But in practice we find that reusing these hyperparameters works well for a given frozen model and level of abstraction in the task. Finally, the LLM used must both cache and expose intermediate activations at the given layer (Bloom and Nanda 2022), which rules out user control of commercial models (though backend use is still possible).
@@
Activation engineering vs fine-tuning. The first advantage of ActAdd  is simple efficiency: the method requires no backward passes and can thus run on any machine that can perform inference rather than training (which, for frontier models, is millions of times more computationally demanding (Fuller 2022). Implementation effort is also greatly reduced; only forward passes are required to find a suitable ($p_{+},p_{-})$ and no labelled data is required (let alone the hundreds of examples of normal finetuning). We discovered most of the example contrast pairs in Appendix Table 6 in minutes. Together, these mean that even nontechnical users can benefit from rapid feedback with roughly the same difficulty as hand-crafted prompt engineering. See Appendix D for more discussion.
@@
Activation engineering vs prompt engineering. Activation additions can be continuously weighted, while prompts are discrete (since a token is either present, or not). To more intensely steer the model to generate wedding-related text, our method does not require any edit to the prompt, but instead just increasing the injection coefficient. See Appendix B for suggestive experiments on ActAdd vs prompting. Unlike the use of extensive ‘meta-prompts’ (prepended natural language specifications), activation additions do not take up token space in the model’s context window, and thus reduce another kind of steering overhead. We argue that activation additions generalize prompt engineering (by allowing weights on token embeddings). We hypothesize that prompting works by activating some goals preferentially (by conditioning the model on one part of activation space). But activation additions may allow control over some properties inaccessible to prompts (Appendix B).
Interpretability of LLMs. In most programs, adding values to imprecisely-targeted intermediate memory locations would not yield sensible results. Why expect this from Transformers? A growing consensus is that the activation space of an LLM contains directions which represent high-level latents causally involved in what is generated (Burns et al. 2022; Moschella et al. 2023; Li et al. 2023a; Nanda 2023; Li et al. 2023b). Our hypothesis, following Elhage et al. 2022, is more specific: that neural networks represent features of the input as directions in activation space, that is, with a linear representation (Park et al. 2023). Moreover, the direction in activation space that corresponds to (say) a love-hate latent variable stays approximately the same across a broad class of inputs.Alain and Bengio 2018 use linear probes on residual streams to infer that LLM representations are at least partially linear; if a linear probe can predict some feature of text output from the residuals with high accuracy, this forms evidence that the feature is represented linearly (i.e. as a simple direction) (Nanda 2023). The success of activation addition gives stronger, experimental evidence of feature linearity, demonstrating that models use feature-related information. Consider the central Love - Hate vector example: we add it to the forward pass and so increase love-related completions. This implicit direction is thus causally responsible for steering the rest of the model to love-related completions. This echoes prior (Nanda 2023; Merullo et al. 2023) and contemporaneous work (Park et al. 2023).
@@
Value alignment of LLMs. Activation engineering is a promising way to control LLMs. Successor methods may be able to provide general steering methods (e.g. through some analogue of a Be helpful vector). Alongside contemporaneous work (Li et al. 2023b; Liu et al. 2023), our experiments suggest that activation engineering can flexibly retarget LLM behavior without damaging general performance. We speculate that this involves changing the model’s currently-active (mixture of) goals. Suitably developed, the activation engineering approach could enable safety progress while incurring a very low ‘alignment tax’ – i.e., cost of opting for safety (Askell et al. 2021; Ouyang et al. 2022).
@@ 6 Conclusion
Activation addition complements existing methods and offers users a novel way to interact with LLMs. We demonstrated natural-language control of output sentiment, topic, and style, and that overall factual performance is preserved. We showed that LLMs afford algebraic combination of forward passes and that ActAdd  scales well with model size. Relative to prompt engineering and finetuning, activation engineering has promising properties: it allows us to compose and reweight model goals at inference time, free up context window space, and allow fast feedback at low computational cost. Our results provide evidence about the computational structure of LLM representations.
Acknowledgments and Disclosure of Funding
We thank Peli Grietzer for providing an independent hyperparameter tuning run. We thank Jan Brauner, Andis Draguns, Sören Mindermann and Raymond Douglas for helpful comments on the draft, as well as Andrew Critch, AI_WAIFU, Aryan Bhatt, Chris Olah, Ian McKenzie, janus, Julian Schulz, Justis Mills, Lawrence Chan, Leo Gao, Neel Nanda, Oliver Habryka, Olivia Jimenez, Paul Christiano, Peter Barnett, Quintin Pope, Tamera Lanham, Thomas Kwa, and Tristan Hume for comments on an earlier draft. We thank Rusheb Shah for engineering assistance. We thank Garrett Baker for running tests on GPT-J (6B). We thank an anonymous ICML reviewer for their extremely thoughtful comments.
6.1 Contributions
@@
MacDiarmid: most of the code, experiments, and quantitative results.
@@
Vazquez: wrote part of text, experiments on LLaMA-3, OPT, GPT-2.

### v5

@@ 4 Results: Activation Addition works
A summary of all experiments can be found in Table 5.333Code repository for our experiments: https://zenodo.org/records/13879423.
4.1 ActAdd intuitively modifies next-token probabilities
@@
We consider the OpenWebText corpus (Peterson et al. 2018). Our running example is the “wedding” topic vector produced by setting $p_{+}=$ weddings, $p_{-}=$   ’, $l=16$, $c=1$.
4.1.1 ActAdd reduces perplexity on a target topic
Figure 2: The perplexity ratio compares the relative predictive performance of ActAdd and an unmodified model. Lower is better. Adding the wedding steering vector improves performance on wedding-related text while preserving performance on unrelated text.
For each document $d_{i}\in D$ in OpenWebText (Peterson et al. 2018), we first calculate the frequency of wedding-related words.444wedding, weddings, wed, marry, married, marriage, bride, groom, and honeymoon. If a document contains one of these words, the document is considered wedding-related. We randomly sample 300k documents, half of which are wedding-related.
We split the documents into sentences and measure GPT-2-XL’s perplexity on both the wedding-related and wedding-unrelated sentences. If the model is effectively steered to generate wedding-related text, it should assign that text higher probability (and thus achieve lower perplexity). For more details, see Section C.1.
@@ 4.1.1 ActAdd reduces perplexity on a target topic
Figure 2 shows the ActAdd perplexity relative to the unmodified model. In sentences where the injected topic (weddings) is more relevant, ActAdd’s perplexity is lower and predictive performance increases.
4.1.2 ActAdd’s impact on token probabilities
To test if the intervention is affecting relevant tokens or reducing perplexity in some spurious way, we observe the shift in the distribution of token log probabilities. We do this by randomly sampling 500 documents from the above OpenWebText sample and recording the log-probabilities assigned by the baseline and steered models. This results in a dataset of about 500k tokens, of which 29k are unique. We then group by token, filter for tokens with $>\!\!20$ instances in the dataset, and calculate the mean perplexity difference between the ActAdd and baseline models. By displaying these as a Q-Q plot (Gnanadesikan & Wilk 1968), we can inspect outlier shifts in token probability.
@@ 4.1.2 ActAdd’s impact on token probabilities
Appendix Figure 8 shows the resulting mean log-probability difference distribution. We see that is approximately normal for the bulk of the tokens but with clearly heavy tails. The positive tail is significantly heavier than the negative tail, suggesting that one set of tokens are reliably increased in probability, with a smaller set of tokens reliably decreased to a lesser extent. Outlier tokens can be found in Appendix Table 11. The probabilities most increased on average are primarily wedding-related. The bottom tokens share no obvious theme and show a significantly lower absolute change in probability.
4.1.3 ActAdd steers the model to discuss weddings
At what layer are steering vectors most effective? Sweeping over GPT-2-XL injection layers for the wedding vector, we measure the average count of wedding-related words given a steering vector injected at each layer.
Figure 3: P(steered completion contains wedding-related words) as a function of injection layer.
@@ 4.1.3 ActAdd steers the model to discuss weddings
The intervention is already effective at the very first layer, rises in effectiveness until layer 6, and then declines. For the optimal injection site, we see $>$90% success in topic steering (compared to a $\sim$2baseline). Figure 3 shows the results of the layer sweep.
4.2 ActAdd can control what the model talks about
Method. Steering vectors can elicit generations on a range of topics – not just weddings. Starting from a generic prompt, we use GPT-3.5 to score whether the generations are about a target topic. Specifically, we generate 100 completions for the unmodified model and 100 for each target single-token ActAdd intervention (each token is about a different topic). Compared to the baseline generations, we record how much more frequently the steered model discusses the target topic.
Results. Figure 4 records a large boost in relevance (5-20%) on all topics at injection coefficient $c=2$ (with the exception of “art”).
@@ 4.2 ActAdd can control what the model talks about
4.3 ActAdd can reduce toxicity
Method. We benchmark toxicity reduction by generating steered continuations from RealToxicityPrompts (Gehman et al., 2020). Following Pei et al. 2023 we use a random subset $n=1,000$. We repeat this sampling 5 times to obtain $p$-values ($t$-test against SOTA), bolding rows which are better with $p<0.05$. For each continuation, we use the Perspective API to score toxicity.
Results. To establish a common scale, we reused the baselines and PREADD results from Pei et al. 2023, adding Air-Decoding Zhong et al. 2023 and FUDGE Yang & Klein 2021. This yields 6 baselines to compare ActAdd against. (We also considered Gu et al. 2022 (which reported 0.043 toxicity), but we could not reproduce the results; also, their disfluency (54.6) is too high for practical use.) We compare to ActAdd using OPT (Zhang et al. 2022b) and LLaMA-3 (Meta 2024). [^5] We do not compare against finetuning because we wish to consider lighter-weight interventions which require minimal gradient updates.
As shown in Table 3, ActAdd-OPT has 8% lower toxicity than the second-best, PREADD-D-OPT, and ActAdd-LLaMA-3 gives a 5% drop over LLaMA-3 with a very small fluency penalty.
Table 3: Results on RealToxicityPrompts (random n=1000). The OPT used is 6.7B parameters, LLaMA-3-8B. Bold is $p<0.05$ against second-best. Gray text denotes numbers reported by Pei et al. 2023 (PREADD), Yang & Klein 2021 (FUDGE), or Zhong et al. 2023 (Air-Decoding). More recent models are less toxic by default. However, ActAdd-OPT is the least toxic of the OPT interventions and even outperforms an unsteered LLaMA-3.
Control Type Method Model Toxicity $\downarrow$ (Dis)Fluency $\downarrow$ Relevance $\uparrow$
@@ 4.3 ActAdd can reduce toxicity
Prompting baseline OPT .200 54.3 .294
Steering vector ActAdd OPT .112 13.8 .329
Controlled gen. FUDGE GPT-2-M .128 22.1 .329
Contrast. decoding PREADD-S OPT .134 51.7 .290
Contrast. decoding PREADD-D OPT .122 56.6 .326
Gradient-guided gen. Air-Decoding GPT-2-L .185 48.3 -
@@
Steering vector ActAdd LLaMA3 .108 6.7 .365
4.4 ActAdd can control sentiment
Method. To evaluate sentiment, we use the Stanford IMDb dataset (Maas et al., 2011). Our goal is for the model to continue each review but with the opposite sentiment. We compute the proportion of generated outputs with the desired sentiment, as classified by a model finetuned on sentiment data, SiEBERT (Hartmann et al. 2023). For quality controls, we follow the conventional use of conditional perplexity to mark (dis)fluency, obtained using GPT-3 davinci-002 logprobs. We use cosine similarity between the prompt and continuation sentence embeddings to gauge the relevance of text in $[0,1]$. We evaluate sentiment changes from positive to negative and vice versa on a random subset of $n=1,000$ and repeat to obtain $p$-values.
Table 4: Results on IMDb sentiment. “Steering” denotes the probability of changing sentiment classification (called “success” in the baselines’ papers). Bold results represent $p<0.05$ compared to the second-best. Gray text denotes numbers reported by Pei et al. 2023. Underline denotes best steered result. Fluency is worse under all steering methods; 1.5x to 3x worse for ActAdd, 7x worse for PREADD.
@@ 4.4 ActAdd can control sentiment
Method Steering $\uparrow$ Disfluency $\downarrow$ Relevance $\uparrow$ Steer. $\uparrow$ Disflu. $\downarrow$ Rel. $\uparrow$
ActAdd-OPT 0.432 24.2 0.387 0.564 20.95 0.363
ActAdd-LLaMA3 0.268 8.6 0.354 0.669 15.2 0.275
OPT-Baseline 0.175 8.95 0.430 0.445 9.38 0.423
LLaMA3-Baseline 0.138 5.8 0.437 0.417 6.09 0.426
OPT-Prompt 0.307 53.5 0.298 0.365 50.9 0.287
FUDGE 0.532 25.1 0.311 0.551 22.7 0.320
PREADD-S-OPT 0.631 68.4 0.253 0.624 67.1 0.258
Results. Table 4 shows that our method is competitive on a conventional measure of sentiment control (Maas et al. 2011). We obtain state of the art success at steering from negative to positive sentiment. While. The only method which outperforms ActAdd in the positive to negative direction incurs a large penalty to fluency (68.4 vs 24.2, when matching methods on the same pretrained model) and relevance.
4.5 ActAdd preserves the model’s general knowledge
Method. We use ConceptNet from the LAMA benchmark, a general knowledge dataset (Petroni et al. 2019, $n=29,774$ sentences, see Appendix Table 10). The model is given a prompt and then has to predict a factual completion. The task is intended for both causal and masked models, so some examples are difficult for causal-attention models (like GPT-2) due to the extremely limited context.
For each sentence, we run the model on its prompt with and without the wedding activation addition. $P@K$ is the probability that the expected label is among the model’s top-$K$ predicted tokens, conditioned on the prompt. We score the baseline and modified models by calculating mean $P@K$ values for a range of $K$. Finally we plot these for both modified and unmodified models over a range of $K$ values.
Results. Figure 5 shows that on the ConceptNet benchmark of factual questions, our method has a negligible impact on off-target answer probabilities.
@@ 5 Discussion
Algebraic combination of forward passes
ActAdd can be viewed as composition of separate forward passes. For example, we compose $\mathbf{h}_{+}$, $\mathbf{h}_{-}$ and $\mathbf{h}^{*}$ to produce steered output. We were surprised that forward passes can “compose” in this way, despite the model not being trained to allow this operation. The composability of forward passes is itself evidence for compositional representations (Olah 2023), independent of the evidence from task-composition arithmetic on weights (Ilharco et al. 2023).
@@
To steer the model using an ActAdd vector, the user supplies the injection coefficient $c$ and the intervention layer $l$. So far we have had success with fixing the sequence alignment $a=1$. Overall, these free hyperparameters make ActAdd less user-friendly than simple prompt engineering. Thankfully, the user does not have to perform a fresh hyperparameter sweep for each use case; in practice, intervention hyperparameters are stable. We include examples of failed steering vectors in Appendix Table 7. We also have not examined ActAdd’s potential impact on reasoning. ActAdd is not immediately applicable given only API access to a model. The model must both cache and expose intermediate activations at the given layer (Bloom & Nanda 2022). Currently, APIs generally do not allow for this.
Activation engineering vs finetuning
Finetuning is better understood and more flexible – we doubt that activation engineering can e.g. teach a model a new skill. However, finetuning is significantly more costly and may not be able to elicit the same kinds of capabilities which activation engineering can elicit.
@@
The first advantage of ActAdd is efficiency: the method requires no backward passes and can thus run on any machine that can perform inference rather than training. Implementation effort is also greatly reduced; only forward passes are required to find a suitable ($p_{+},p_{-})$ and minimal labelled data is required - just the steering prompt pair. We discovered most of the example contrast pairs in Appendix Table 6 in minutes. All things considered, even nontechnical users can benefit from rapid feedback and relatively easy iteration
Activation engineering vs prompt engineering
@@
Activation additions can be continuously weighted, while prompts are discrete – a token is either present, or not. To more intensely steer the model to generate wedding-related text, our method does not require any edit to the prompt, but instead just increasing the injection coefficient. See Appendix B for suggestive experiments on ActAdd vs prompting. Unlike system prompts, activation additions do not take up token space in the model’s context window, although this is a small benefit in the era of multi-million token context windows.
While prompting is more flexible and even cheaper than ActAdd, activation additions may elicit capabilities which prompting cannot (as evidenced by our superior results over prompting; see also the speculation in Section 1).
Interpretability of LLMs
In most programs, adding values to imprecisely targeted intermediate memory locations would not yield sensible results. Why expect this from Transformers?
A growing consensus is that the activation space of an LLM contains directions which represent high-level latents causally involved in what is generated (Burns et al. 2022; Moschella et al. 2023; Li et al. 2023a; Nanda 2023; Li et al. 2023b).
Our hypothesis, following Elhage et al. 2022, is more specific: that neural networks represent features of the input as directions in activation space, that is, with a linear representation (Park et al. 2023). Moreover, the direction in activation space that corresponds to (say) a love-hate latent variable stays approximately the same across a broad class of inputs.
Alain & Bengio 2018 use linear probes on residual streams to infer that LLM representations are at least partially linear; if a linear probe can predict some feature of text output from the residuals with high accuracy, this forms evidence that the feature is represented linearly (i.e. as a simple direction) (Nanda 2023).
The success of activation addition gives stronger, experimental evidence of feature linearity, demonstrating that models use feature-related information. Consider the central Love - Hate vector example: we add it to the forward pass and so increase love-related completions. On the examined prompts, this direction is responsible for steering the rest of the model towards love-related completions. In general, steering vectors establish causality, at least in the limited set of contexts examined.
Value alignment of LLMs
@@
Activation engineering is a promising way to control LLMs. Successor methods may be able to provide general steering methods (e.g. through some analogue of a Be helpful vector). Alongside contemporaneous work (Li et al. 2023b; Liu et al. 2023), our experiments suggest that activation engineering can flexibly retarget LLM behavior without damaging general performance. We speculate that ActAdd changes the model’s currently active mixture of goals and priorities. Suitably developed, the activation engineering approach could enable safety progress while preserving overall capabilities
@@ 6 Conclusion
While methods like prompt engineering, controlled decoding, and finetuning have benefits, they fail to elicit full capabilities from language models. To more reliably elicit these abilities, activation engineering strategically perturbs activations at inference time. In particular, we introduced Activation Addition to steer models by shifting their inference-time activations along a certain direction (like the “Love”-“Hate” vector). ActAdd is lightweight and effective, achieving SOTA on toxicity reduction and sentiment shift while retaining overall model capabilities. ActAdd demonstrates the promise of activation engineering. We look forward to future work realizing this promise.
Reproducibility Statement
Our code is available here: https://zenodo.org/records/13879423. The following is an exhaustive list of models used, sampling strategies used, and searches run:
Data processing
To curate a wedding-related subset of OpenWebText, we retained documents with wedding-related words (see Section 4.1.1). The only pre-processing performed is to remove sequences of null characters. Each document is split into sentences $s_{j}\in d_{i}$ using the Punkt tokenizer (Strunk 2013).
Models
After observing success with GPT-2-XL, to replicate our results, we subsequently repeated the same experiments with Llama-1-13B (Touvron et al. 2023) and GPT-J-6B (Wang & Komatsuzaki 2021). Our toxicity and sentiment experiments use OPT (Zhang et al. 2022b) and LLaMA-3-8B Meta 2024. See Appendix E for details. We use all-MiniLM-L6-v2 (Reimers & Gurevych 2019) to compute sentence embeddings to calculate relevance using cosine similarity. For the success score, we use the SiEBERT (Hartmann et al. 2023) sentiment classifier. We perform sentiment classificaton with the SiEBERT classifier (Hartmann et al., 2023).
APIs
For scoring toxicity, we use https://www.perspectiveapi.com/. For scoring fluency, we use OpenAI davinci-002. The PREADD baseline instead used the discontinued davinci-001 model.
Seed
Sampling hyperparameters
We precommitted to fixed sampling hyperparameters, selected before experiments began. We held them fixed throughout our data collection. Those sampling hyperparameters were temperature$=1.0$, freq_penalty$=1.0$, and top_p=$0.3$. Since this top_p value seemed a bit unusual to us in retrospect, we invited an external researcher to reproduce this process with an unmodified GPT-2-XL and report the best sampling hyperparameters they found. This second experiment was blinded, as they did not know the values we used. They found that temperature$=0.6$ and top_p$=0.5$ produced better GPT-2-XL capabilities. We reran all our qualitative results at this setting, and they all reproduced (subjectively, more impressively).
In replicating the unsteered OPT sentiment baseline, we find that the NegToPos direction is consistently higher success than PosToNeg. This holds across different combinations of model hyperparameters, including those in Pei et al. 2023. However, PREADD Pei et al., 2023 reports similar success results for both (i.e. a much lower NegToPos success). The OPT results use our calculated values.
Reporting the best of $K$ completions
We generated $K=3$ completions for each qualitative demonstration, for both normal and steered forward-passes. Appendix Table 6, shows the subjectively most compelling completion pair out of the first three seed-00 completion-pairs. You can see all top-3 completions for the entries in this notebook: tinyurl.com/actadd3. We share activation additions which work well. We iterated over contrast pairs to get these to work, although several striking results were generated within [first author’s] first hour of using the technique. Out of the 12 activation additions we thought demonstrated a distinct ability of the method, we decided not to include 1 because its first three seed-00 completions were unusually unimpressive. We include the remaining 11 in Table 6.
ActAdd  hyperparameters ($l,c$)
Hardware:
GPU: Nvidia RTX A5000, CPU: Intel Core i9-10900X CPU @ 3.70GHz. 24GB GPU RAM, 32GB system RAM
Relevant libraries and frameworks:
Operating system: Ubuntu 22.04.1 LTS, numpy: 1.24.3, pandas: 2.0.1, torch: 1.13.1, transformer-lens: 1.4.0.
Author Contributions
@@
MacDiarmid: most of the main library code.
@@
Vazquez: wrote part of text, conducted toxicity, sentiment control, and other experiments on LLaMA-3, OPT, GPT-2.
Acknowledgments
We thank Peli Grietzer for providing an independent hyperparameter tuning run. We thank Alex Cloud, Jan Brauner, Andis Draguns, Sören Mindermann and Raymond Douglas for helpful comments on the draft, as well as Andrew Critch, Aryan Bhatt, Chris Olah, Ian McKenzie, janus, Julian Schulz, Justis Mills, Lawrence Chan, Leo Gao, Neel Nanda, Oliver Habryka, Olivia Jimenez, Paul Christiano, Peter Barnett, Quintin Pope, Tamera Lanham, Thomas Kwa, and Tristan Hume for comments on an earlier draft. We thank Rusheb Shah for engineering assistance. We thank Garrett Baker for running tests on GPT-J (6B) We thank an anonymous ICML reviewer for their extremely thoughtful comments.